Calculate the mole fraction of ethanol and water in a sample of rectified spirit which contains 92% ethanol by mass. Please give me answer

Grade:12

3 Answers

Gowtham
6 years ago
the question states that the solution has 92% ethanol by mass
so if the solutions weighs 100gms , it contains 92gms of ethanol and 8gms of water 
no.of moles of ethanol = weight of ethanol / molecular weight of ethanol
molecular weight of ethanol is 46
no. of moles of ethanol is 92/46 i.e 2
no.of moles of water is 8/18 i.e 0.222
 
mole fraction of ethanol = no.of moles of ethanol / total number of moles
                                      = 2/2.222
                                      =0.90 (approx)
mole fraction of water =1-mole fraction of ethanol
                                    =1-0.9
                                    =0.1 (approx)
ankit singh
3 years ago
Rectified spirit which contains 46% ethanol by mass.
In 100 g spirit has 46 grams of ethanol and 54 grams of water.
Rectified spirit which contains 46% ethanol by mass.
In 100 g spirit has 46 grams of ethanol and 54 grams of water.
Moles of ethanol=n_e=\frac{46 g}{46 g/mol}=1 mol
Moles of water n_w=\frac{54 g}{18 g/mol}= 3 mol
\chi_e=\frac{n_e}{n_e+n_w}=\frac{1}{1+3}=0.25
\chi_w=\frac{n_w}{n_e+n_w}=\frac{3}{1+3}=0.75
Mole fraction of ethanol is 0.25.
Mole fraction of water is 0.1
Aryan
one year ago
 =  
2
 H 
5
 OH
 +n 
2
 O
 
2
 H 
5
 OH
 
 
 
Molar mass of ethanol = 46g/mol
 
Molar mass of water = 18g/mol
 
Mass of ethanol =85g
No of moles of ethanol =  
46
85
  = 1 =  
2
 H 
5
 OH
 +n 
2
 O
 
2
 H 
5
 OH
 
 
 
Molar mass of ethanol = 46g/mol
 
Molar mass of water = 18g/mol
 
Mass of ethanol =85g
No of moles of ethanol =  
46
85
  = 1.85mol
 
Mass of water = 100−85 = 15g
 
No of moles of water =  
18
15
  = 0.833mol